# A Zero-Added-Loss Multiplexing (ZALM) Source Simulation

### v3

@@ I Introduction
Robust entanglement distribution is a foundational requirement for the future quantum internet, as laid out by [kimble2008quantum] and [wehner2018quantum]. While much research has focused on optimising the rate and fidelity of this entanglement distribution, the next step is integrating with quantum memories to form functional quantum memory networks. Adopting the principles of systems engineering, we argue that a holistic, end-to-end analysis of the network architecture is critical. This perspective integrates the performance of quantum entanglement sources, the fidelity degradation across quantum channels, and the efficiencies of quantum memories. To address this, we model a novel quantum source specifically engineered for high-efficiency loading into quantum memories, thereby improving the performance of the entire quantum network.
@@
Quantum entanglement sources come in many forms, with different natural wavelengths, emission efficiencies, and bandwidths, as described by [eisaman2011invited]. Placing these sources into two main categories, we can look at their emission properties as either deterministic, such as the very precise wavelengths emitted by quantum dots, or probabilistic, whereby the specific emission wavelength of a generated photon pair is a stochastic variable distributed across a broad spectral bandwidth centred at a known degeneracy point. Although deterministic sources simplify network setups, the broadband capabilities of probabilistic sources allow for greater use of the available spectrum. This is enabled by adapting common classical network management techniques for routing and spectrum allocation, see [jsac_rsa].
@@
To scale quantum networks beyond simple point-to-point links, quantum memories are an essential component. They provide the crucial capability to synchronise operations between distant nodes and to mitigate the inherently probabilistic nature of entanglement generation and swapping protocols. As such, memories form the cornerstone of quantum repeater architectures. The search for a practical physical implementation has led to significant research in solid-state systems. For instance, [weber2010quantum] showed that defects, or vacancies, in diamond are a practical and robust mechanism to provide quantum memories. Further work by [bradac2019quantum] highlighted the attributes of the other group IV elemental defects in diamond, with Silicon showing considerable coherence times, > 1s, and a more efficient spin-photon interface. However, successfully interfacing with these memories is a significant challenge, with [bersin2024telecom] and [PhysRevApplied.22.044013] detailing the specific requirements on the incoming photons’ properties, such as bandwidth and wavelength matching.
@@
The Zero-Added-Loss Multiplexing (ZALM) protocol introduced by [Chen_ZALM] provides for a broadband source with high memory loading efficiency. As there is currently no agreed implementation of the transmitter (Tx) component of this protocol and there are numerous configuration options of each of the transmitter subcomponents, we created a configurable simulation to determine the optimal transmitter architecture for maximising entanglement distribution rates and fidelity.
@@ III-B2 Hong–Ou–Mandel Visibility
No system is ideal, and a major source of decoherence is introduced by the Hong–Ou–Mandel (HOM) visibility of the photons that interact in the beam splitter, see [HOMvisibility]. In the case of the ZALM source, the HOM visibility is concerned with how closely the two photons overlap spectrally, temporally, and in polarisation, with perfect overlapping in all three degrees of freedom producing completely indistinguishable photons. Anything less than perfect reduces the fidelity of the entanglement to the point that fidelity is zero if the photons are completely distinguishable.
@@ III-D1 Delay Line
The singlet state is the preferred choice for the quantum memory, as it can be loaded with very high (unit) fidelity using the augmented Duan-Kimble protocol described by [PhysRevApplied.22.044013].
Lastly, [shapiro2024entanglement] provide a simple circuit that could be used to convert the polarisation encoded qubits into time-bin encoded qubits for more robust transmission over fibre.
@@ IV Implementation / Results
This ZALM source simulation is implemented in NetSquid, developed by [coopmans2021netsquid], and uses the QSI Controller and APIs from [QSI_Library_2022]. It can run standalone in NetSquid, or as part of a larger system within QSI.
@@ IV-C Results
Changing the SPDC_DEGENERACY_BANDWIDTH_FWHM_NM to 1.0 and the DETECTOR_TYPE to PNR improves the ebit rate considerably, whilst the average fidelity remains roughly the same, see Fig. 15.
@@ VI Future Work
In the future work, we will adjust the exiting SPDC parameters to implement new advances in ZALM source research, such as the phase-matched spectral islands architecture introduced by [zalm-islands].
@@ Acknowledgments
The data is available from the authors upon reasonable request. Code is freely available on GitHub, published by [Horgan_ZALM_Source_Simulation_2025], under the GPLv3 license.

### v4

@@ I Introduction
Robust entanglement distribution is a foundational requirement for the future quantum internet [kimble2008quantum, wehner2018quantum]. While much research has focused on optimising the rate and fidelity of this entanglement distribution, the next step is integrating with quantum memories to form functional quantum memory networks. Adopting the principles of systems engineering, we employ a holistic, end-to-end analysis of the network architecture, following a modular approach consistent with prominent quantum network simulation frameworks [Wu_2021, diadamo2020qunetsim]. This perspective integrates the performance of quantum entanglement sources, the fidelity degradation across quantum channels, and the efficiencies of quantum memories. While recent studies have successfully modelled entanglement generation between absorptive memories [9951205] and simulated atomic ensembles [zhou2023simulator], these works primarily focus on the link-layer physics. We expand upon this body of knowledge by bridging the gap between component-level quantum optics and system-level network metrics. To address this, we model a novel quantum source specifically engineered for high-efficiency loading into quantum memories, demonstrating how optimizing the source-memory interface is critical for improving the performance of the entire quantum network.
@@
Quantum entanglement sources come in many forms, with different natural wavelengths, emission efficiencies, and bandwidths [eisaman2011invited]. Placing these sources into two main categories, we can look at their emission properties as either deterministic, such as the very precise wavelengths emitted by quantum dots, or probabilistic, whereby the specific emission wavelength of a generated photon pair is a stochastic variable distributed across a broad spectral bandwidth centred at a known degeneracy point. Although deterministic sources simplify network setups, the broadband capabilities of probabilistic sources allow for greater use of the available spectrum. This is enabled by adapting common classical network management techniques for routing and spectrum allocation [jsac_rsa].
@@
To scale quantum networks beyond simple point-to-point links, quantum memories are an essential component. They provide the crucial capability to synchronise operations between distant nodes and to mitigate the inherently probabilistic nature of entanglement generation and swapping protocols. As such, memories form the cornerstone of quantum repeater architectures. The search for a practical physical implementation has led to significant research in solid-state systems. For instance, defects, or vacancies, in diamond are a practical and robust mechanism to provide quantum memories [weber2010quantum]. Further work highlighted the attributes of the other group IV elemental defects in diamond, with Silicon showing considerable coherence times, greater than 1 s, and a more efficient spin-photon interface [bradac2019quantum]. However, successfully interfacing with these memories is a significant challenge, with specific requirements on the incoming photons’ properties, such as bandwidth and wavelength matching [bersin2024telecom, PhysRevApplied.22.044013].
@@
The Zero-Added-Loss Multiplexing (ZALM) protocol provides for a broadband source with high memory loading efficiency [Chen_ZALM, shapiro2024entanglement]. As there is currently no agreed implementation of the transmitter (Tx) component of this protocol and there are numerous configuration options of each of the transmitter subcomponents, we created a configurable simulation to determine the optimal transmitter architecture for maximising entanglement distribution rates and fidelity.
@@ III-B2 Hong–Ou–Mandel Visibility
No system is ideal, and a major source of decoherence is introduced by the Hong–Ou–Mandel (HOM) visibility of the photons that interact in the beam splitter [HOMvisibility]. In the case of the ZALM source, the HOM visibility is concerned with how closely the two photons overlap spectrally, temporally, and in polarisation, with perfect overlapping in all three degrees of freedom producing completely indistinguishable photons. Anything less than perfect reduces the fidelity of the entanglement to the point that fidelity is zero if the photons are completely distinguishable.
@@ III-D1 Delay Line
The singlet state is the preferred choice for the quantum memory, as it can be loaded with very high (unit) fidelity using the augmented Duan-Kimble protocol [PhysRevApplied.22.044013].
Lastly, a simple circuit could be used to convert the polarisation encoded qubits into time-bin encoded qubits for more robust transmission over fibre [shapiro2024entanglement].
@@ IV Implementation / Results
This ZALM source simulation is implemented in NetSquid [coopmans2021netsquid], and uses the QSI Controller and APIs [QSI_Library_2022]. It can run standalone in NetSquid, or as part of a larger system within QSI.
@@ IV-C Results
Changing the SPDC_DEGENERACY_BANDWIDTH_FW-HM_NM to 1.0 and the DETECTOR_TYPE to PNR improves the ebit rate considerably, whilst the average fidelity remains roughly the same, see Fig. 15.
@@ VI Future Work
In the future work, we will adjust the exiting SPDC parameters to implement new advances in ZALM source research, such as the phase-matched spectral islands architecture [zalm-islands].
@@ Acknowledgments
The data is available from the authors upon reasonable request. Code is freely available on GitHub [Horgan_ZALM_Source_Simulation_2025], under the GPLv3 license.